OpenStudy (anonymous):

Approximate the integral using Simpson's rule S10 integral from 1 to 2 lnx^(3/2) dx

OpenStudy (anonymous):

Pigg, are you going to answer this, or give me 3 mins for me to figure it out?

OpenStudy (anonymous):

All you do is just plug the information given into Simpson's rule ( http://tutorial.math.lamar.edu/Classes/CalcII/ApproximatingDefIntegrals_files/eq0026MP.gif) a=1, b=2, n=10, thus change in x:[(b-a)/n]=1/10, more info at http://tutorial.math.lamar.edu/Classes/CalcII/ApproximatingDefIntegrals.aspx

OpenStudy (anonymous):

\[\Delta x = (2-1)/n = 1/10.\] so do EACH one of the subdivisions separately.... f(a) = ln(1)^(3/2) = (3/2)*ln(1) f(a+\[\Delta x\] = ln(1+1/10)^(3/2) = (3/2)*ln(1.1) and so on... you got the busy work, and then sum all 10 divisions up. or you can cheat and use the calculator Flyinpigg linked.

OpenStudy (anonymous):

try clicking on the links pal

OpenStudy (anonymous):

Thanks alot for the help, it cleared up some confusion. : )